# Student member, IEEE

### v1

@@ 4.3 Results and analysis
Fig. 9 displays the $C_{P}$ variations of multiple wing sections with span width of 28.8%, 44.9%, 54.3%, 68.1%, 75.1%, and 81.8% at $AoA$ of 20°. In subgraphs (a) and (b), the black curves closely align with the red points in the lower part of each curve, indicating that our method effectively captures the $C_{P}$ variations on the upper surface of the DLR-F11 wing. In subgraphs (c) and (d), all curves exhibit similar variations. In subgraphs (e) and (f), the black curves at the first and third wing stage closely match the red points. The black lower curve of the second wing stage are closer to red points than other curves. In addition, only one or two points in the black upper curve are inaccurate, and the remaining points are accurate. In general, our method accurately predicts $C_{P}$ variations on the first and third wing stage, as well as on the upper surface of the second wing stage. However, our method are inaccurate in predicting $C_{P}$ value of few points on the lower surface of the second wing stage (Notice: the upper curve in Fig. 9 denotes the $C_{P}$ variations on the lower surface of the DLR-F11 wing, and the lower curve denotes the $C_{P}$ variations on the upper surface of the DLR-F11 wing).
@@
Fig. 10 illustrates $C_{P}$ variations of the same wing sections with the same span widths at $AoA=18.5$°. In subgraphs (a) and (b), the lower part of the blue curves of the third wing stage exhibits a zigzag pattern. In contrast, the black curves in the corresponding region are relatively smooth. This observation indicates that DAN is difficult to accurately predict the $C_{P}$ variation on the upper surface of the third wing stage, while our method can predict more accurate $C_{P}$ variations than DAN in this region. The green curves almost coincide with the black curves. However, comparing the lower parts of the second and third wing stage, it is evident that MLP is not as accurate as our method. In subgraphs (c) to (f), the zigzag pattern of the blue curves are more obvious, especially within the curves corresponding to the third wing stage. These observations suggest that the geometric features proposed in this paper offer a more accurate representation of 3D wings than images (DAN use images of wing sections to predict the $C_{P}$ distributions of the DLR-F11 wing).
5 CONCLUSION
@@
In this paper, we propose a learning method that incorporating Riemannian geometric features for predicting the $C_{P}$ distributions on wing surfaces in 3D space. We demonstrate that the geometry of a wing in 3D space can be described by a set of self-intersection-free Bézier surfaces, and further demonstrate that three geometric features (Riemannian metric, connection and curvature) can be used to represent the geometric structure of a wing in 3D space. The experimental results show that our method which incorporating the geometric features from 9 points (including the additional 8 neighbors) into the multi-feature learning module reduces the average prediction MSE of $C_{P}$ by 8.41% for the DLR-F11 aircraft test set.

### v2

@@ 4.3 Results and analysis
Fig. 9 displays the $C_{P}$ variations of multiple wing sections with span width of 28.8%, 44.9%, 54.3%, 68.1%, 71.5% and 75.1% at $AoA$ of 20°. In subgraphs (a) and (b), the black curves closely align with the red points in the lower part of each curve, indicating that RGFiL_2 effectively captures the $C_{P}$ variations on the upper surface of the DLR-F11 wing. In subgraphs (c) and (d), all curves exhibit similar variations. In subgraphs (e) and (f), the black curves at the first and third wing stage closely align with the red points. The black lower curve of the second wing stage are closer to the red points than other curves. In general, RGFiL_2 accurately predicts $C_{P}$ variations on the three wing stages, especially on the upper surface of the second wing stage. However, RGFiL_2 are inaccurate in predicting $C_{P}$ value of few points on the lower surface of the second wing stage (Notice: the upper curve in Fig. 9 denotes the $C_{P}$ variations on the lower surface of the DLR-F11 wing, and the lower curve denotes the $C_{P}$ variations on the upper surface of the DLR-F11 wing).
@@
Fig. 10 illustrates $C_{P}$ variations of the same wing sections with the same span widths at $AoA=18.5$°. In all subgraphs, the lower part of the blue curves and green curves on the third wing stage deviate from the red points, and the blue curves deviate more significantly. In contrast, the black curves in the corresponding region are closer to the red points. This observation indicates that DAN and MLP_2 are difficult to accurately predict the $C_{P}$ variation on the upper surface of the third wing stage, while RGFiL_2 can predict more accurate $C_{P}$ variations than DAN and MLP_2 in this region. In subgraphs (a), (d), (e) and (f), the lower part of the black curves on the second wing stage are closer to the red points than the corresponding part of both the blue curves and green curves. These observations suggest that RGFiL_2 proposed in this paper offer a more accurate representation of 3D wings than DAN and MLP_2.
Figure 11: The weights of neural representations for three random points with $AoA=18.5$°.
Figure 12: The weights of neural representations for three random points with $AoA=20$°.
4.4 Analysis: The Contribution of Riemannian Geometric Features
In this section, we analyze the importance of five neural representations during the training process of RGFiL_2. Figure. 11 shows the weight variations of neural representations (i.e., $f_{1}(\vec{\mathbf{x}}_{1})\sim f_{5}(\vec{\mathbf{x}}_{5})$) for three random points $P_{1}$, $P_{2}$ and $P_{3}$ on the DLR-F11 manifold with $AoA=18.5$°. Subgraphs (a) to (c) show the weight variations of neural representations on $P_{1}$, $P_{2}$ and $P_{3}$, respectively. Subgraph (d) shows the position of $P_{1}$, $P_{2}$ and $P_{3}$ on the DLR-F11 manifold. When epoch $\in[1000,2000]$, we see that 1) the weight of $f_{2}(\vec{\mathbf{x}}_{2})$ decreases, while the weights of both $f_{4}(\vec{\mathbf{x}}_{4})$ and $f_{5}(\vec{\mathbf{x}}_{5})$ increase; 2) the weight of $f_{1}(\vec{\mathbf{x}}_{1})$ does not significantly change; 3) the weights of $f_{3}(\vec{\mathbf{x}}_{3})$ decreases.
Figure. 12 shows the weight variations of $f_{1}(\vec{\mathbf{x}}_{1})\sim f_{5}(\vec{\mathbf{x}}_{5})$ for $P_{1}$, $P_{2}$ and $P_{3}$ on the DLR-F11 manifold with $AoA=20$°. We see that 1) the weight of $f_{2}(\vec{\mathbf{x}}_{2})$ is biggest; 2) the weights of $f_{4}(\vec{\mathbf{x}}_{4})$ and $f_{5}(\vec{\mathbf{x}}_{5})$ are always bigger than those of $f_{1}(\vec{\mathbf{x}}_{1})$ and $f_{3}(\vec{\mathbf{x}}_{3})$; 3) the weights of $f_{1}(\vec{\mathbf{x}}_{1})$, $f_{2}(\vec{\mathbf{x}}_{2})$, $f_{4}(\vec{\mathbf{x}}_{4})$ and $f_{5}(\vec{\mathbf{x}}_{5})$ tend to be stable as the number of iterations increases, but the weight of $f_{3}(\vec{\mathbf{x}}_{3})$ decreases as the number of iterations increases.
Theoretically, the Riemannian metric $g_{ij}$ is an inner product on the tangent space of a manifold. The $g_{ij}$ forms the basis for other Riemannian geometric features, for examples, the distance, angle, connection, and curvature [14]. The connection coefficient $\Gamma_{ij}^{k}$ is instrumental in defining covariant derivatives and parallel transport [40], establishing relationships between tangent spaces at different neighboring points on the aircraft manifold. Connections between different tangent spaces play a crucial role in representing the intrinsic geometric integrity of the aircraft manifold. Scalar curvature $S$ is employed to reflect the bending and deformation properties of the wing surface [41]. The Riemannian geometric features are used to represent the intrinsic geometric properties of the 3D DLR-F11 aircraft, whcih explains why the RGFiL shows better accuracy than other methods. Notably, the geometric meanings of $\Gamma_{ij}^{k}$ and $S$ are more clear and suitable for aircrafts than that of $g_{ij}$, thus the RGFiL keeps reducing the weight of $f_{3}(\vec{\mathbf{x}}_{3})$ during the training process.
5 Conclusion and Future Work
@@
In this paper, we propose RGFiL that incorporating Riemannian geometric features from an arbitrary point and its 8 neighbors to predict the $C_{P}$ distributions on a 3D aircraft. We demonstrate that the geometry of a 3D aircraft can be described by a piecewise smooth manifold, and further demonstrate that three Riemannian geometric features (Riemannian metric, connection and curvature) can be used to represent the geometric integrity of the 3D aircraft. Experimental results show that RGFiL reduces the average prediction MSE of $C_{P}$ by 15.00% for the DLR-F11 aircraft test set.
In the future, we will build stronger relations between Riemannian geometry features and deep learning, for example the connection can be used to control the parallel transport of velocity vectors on aircraft manifold, which can be used to explore the internal mechanism of flow field variations.